OpenStudy (anonymous):

a Goldsmith designed a trapezoidal pendant with the perpendicular distance between AB and CD is 6 centimeters and the area of the pendant ABCD is 36 square centimeters find the length of CD.

OpenStudy (anonymous):

Please help me!!! Let me see the work that you did step by step.

OpenStudy (anonymous):

What is the area of a trapezoid

OpenStudy (anonymous):

A=a+b/2×h

OpenStudy (anonymous):

So how do u think u can find the height

OpenStudy (anonymous):

I don't know that's why I'm asking for help. I really have no ide!!

OpenStudy (anonymous):

Hold on do uu have a photo of this question

OpenStudy (anonymous):

No, but I think I figured it out. Is the answer 4?

OpenStudy (anonymous):

How did u get four

OpenStudy (anonymous):

Thank you for your help!!!😆

OpenStudy (anonymous):

I put "x" for base one and just worked it out

OpenStudy (anonymous):

Here: A=1/2×h (base one + base two) A= 1/2×6 (x+8) A=3(x+8) A= 3×8 + 3× x A= 3x + 24 = 36 - 24. - 24 A= 3x = 12 Divide 3 on both sides and it ends up leaving you with X= 4 after that I just checked my work by plugging it into the formula.